OpenStudy (anonymous):

integral 1/(y^2 + 5y + 6) dy Is there another way to find this other than completing the square on the denominator?

ganeshie8 (ganeshie8):

whats wrong with partial fractions ?

OpenStudy (cwrw238):

this will factor

OpenStudy (anonymous):

i was thinking maybe that but wasn't sure if that would work., also was thinking factor then IBP?

OpenStudy (ribhu):

\[\frac{ 1 }{ (y+2)(y+3) }\]

OpenStudy (ribhu):

\[\frac{ 1 }{ y+2 } - \frac{ 1 }{ y+3 }\]

OpenStudy (cwrw238):

yes

OpenStudy (ribhu):

then separately it can be easily integrated.

OpenStudy (anonymous):

oh wow I didn't know you could do that. ty :)

OpenStudy (ribhu):

thanks @ganeshie8

OpenStudy (anonymous):

How do you go from the factors to the sum? is that by PFD?

OpenStudy (cwrw238):

PFD yes

OpenStudy (anonymous):

ah, ty @ganeshie8, @cwrw238, @ribhu

OpenStudy (anonymous):

im going to try it out now, ty!

OpenStudy (ribhu):

in your inbox

OpenStudy (cwrw238):

thats the best way - learn by practicing yourself

ganeshie8 (ganeshie8):

it helps to remember that EVERY rational function can be integrated using PFD

OpenStudy (anonymous):

for some reason i though you could not do PFD unless you had a binomial in the numerator, thats all I have done it for

OpenStudy (anonymous):

thought*

OpenStudy (anonymous):

ty all, sorry for my ignorance

OpenStudy (cwrw238):

dont be